OpenStudy (anonymous):

Find the Value of x and explain step by step how you got it- Please see below...

OpenStudy (anonymous):



OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

K ill see if i can do it

OpenStudy (anonymous):

k this is what i have 

OpenStudy (anonymous):

correct so far?

OpenStudy (anonymous):

Yep, now multiply it out and solve for x

OpenStudy (anonymous):

What do you mean multiply it out?

OpenStudy (anonymous):

Distribute the 2 on the left and the 3 on the right

OpenStudy (anonymous):

k i have 8x - 2 = 6x + 3. Correct so far?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

well you could subtract 6x from both sides 2x-2=3 then add 2 to both sides 2x=5 and finally divide both sides by 2 x=5/2

OpenStudy (anonymous):

Thanks i know how to do it now